A total of 9 students donated money for a school fundraiser and their average donation was $12. If the average donation for 5 of the students was $16, what was the average donation, for the remaining for students?
A)$5 B)$6 C)$7 D)$8 E)$9

Answer by ikleyn(52867)

.
A total of 9 students donated money for a school fundraiser and their average donation was $12.
If the average donation for 5 of the students was $16, what was the average donation,
for the remaining for students?
A)$5 B)$6 C)$7 D)$8 E)$9
~~~~~~~~~~~~~~~~~~

If the average donation of 9 students was $12, it means that the total amount of donation 
of these 9 students was  9*12 = 108 dollars.


Since the average donation of 5 of the students was $16, their total donation was 5*16 = 80 dollars.


Hence, the total donation of remaining 4 students was  108 - 80 = 28 dollars.


It means that the average donation of these 4 students was  28/4 = 7 dollars.    ANSWER

Solved and explained.